(-2,3) (4,k), m= 1/3 how do you get the answer for the value of k fir the pair of points with the given slope
Answer by stanbon(75887)

(-2,3) (4,k), m= 1/3 how do you get the answer for the value of k fir the pair of points with the given slope
------
Equation:
(k-3)/(4--2) = 1/3
(k-3)/6 = 1/3
---
k-3 = 6/3
k = 5
--------------
Cheers,
Stan H.
------------